# SPECIAL RELATIVITY

### v9

@@ 5.5 A Resynchronization of the Accelerated Clocks
It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame $S_{n}$. The explanation consists in using a clock re-synchronization procedure. Well known that in their original form Maxwell’s equations are valid in the inertial frames. But Maxwell’s equations can be written down only if the Lorentz coordinates has already been specified.
@@
When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction. Using such synchronization procedure we actually select a Lorentz coordinate system for the screen. In this synchronization, we describe the transmission through the aperture using usual Maxwell’s equations. The interval in the accelerated reference system $S_{n}$ will have the diagonal form Eq.(1) for the transmitted light beam.
@@
Figure 16: Aberration of light in the accelerated frame of reference $S_{n}$. The plane wavefront rotates in the accelerated frame after the accelerated clock re-synchronization. The Maxwell’s equations can now be used to describe the transmitted (through the aperture) light beam. According to the Maxwell’s electrodynamics, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
@@
The time $t^{\prime}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation. This time shift has the effect of rotation the plane of simultaneity (that is emitter radiation wavefront) on the angle $-v/c$ in the first order approximation. As a consequence of this, the plane wavefront rotates in the accelerated frame after re-synchronization, Fig. 16. The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the light transmission (through the aperture) description. Then, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light: the two approaches give the same result. The choice between these two different clock synchronizations is a matter of pragmatics. By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear. But we can obtain a more consistent description of these phenomena.
@@
We had to restrict ourselves to finding the diagonalization procedure only for nonrelativistic velocities. The method can be easily generalized for an arbitrary parameter $v/c$. A full diagonalization can be obtained using transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Physical time $t^{\prime}_{n}$ determines the flow of time in a physical process in the accelerated frame $S_{n}$. The $dx^{\prime}_{n}$ is the physical distance between to points of 3-space.
@@ 5.6 Composition of Motions
Until now we have considered only the case in which motion is nonrelativistic in speed. Now we shall be more general and study the case where the motion is at an arbitrary velocity. We already know from our discussion in this chapter that metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the inverse Galilean boost. In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame. Let us consider the transformation of the direction of light propagation. The light (group) velocity transforms as a particle velocity. In this case we use the Galileo addition theorem of velocities for transforming the speed of propagation of light to the accelerated frame. Let us assume that the accelerated frame moves with velocity $v$, then the angle of aberration $\theta_{a}$ is such that   $\tan\theta_{a}=-v/c$. If $v/c\ll 1$ the angle of aberration is small, being approximated by $\theta_{a}=-v/c$.
It is always possible to chose such variables, in which metric of the accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. In fact, substituting $dx_{n}/dt_{n}=-v$ into this equation gives $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. Group velocity of light beam along the fixed stars motion is still given by $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This is modification of the aberration of light law in the Lorentz coordinatization. Suppose that accelerated observer performs a measurement of the light aberration. Since in the Lorentz coordinatization the time and distance have direct physical meaning, the accelerated observer finds that the angle of aberration is such that $\sin\theta_{a}=-v/c$.
@@ 5.7 Discussion
Now let us return to metric diagonalization and try to get better understanding of the relationship between the $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ coordinate systems. We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by accelerated observer? Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and both frames are equivalent. Above we demonstrated that there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. From a mathematical viewpoint , we suggest the following explanation. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$. We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
@@
Where does the asymmetry come from? The electrodynamics equation needs to be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical direction $z$. On the other hand, we know that when we are dealing with Lorentz coordinatization we have effect on the wavefront orientation in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
@@
It is generally accepted that this is purely kinematic effect, involving no forces. Many people will, we think, find this disturbing. They would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation. We discuss the key of this paradox in great detail in the Section 9.1. It is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. Without proof, we may state the results. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the system $S_{n}$ as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process. The rate of a clock depends on the pseudo-gravitational potential at the place where the clock is situated. The time-offset relation can be interpreted as the accomulated time difference between spatially separated clocks. Suppose (for simplicity) that there is a constant acceleration $g$ along $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. The clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks we have $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift has the effect of rotation the plane of simultaneity on the angle $-v/c$.
5.8 A Single Moving Emitter in a Noninertial Frame of Reference
Above in the sections 4.3 - 4.4 we considered a single moving ”plane wave” emitter in the initial inertial frame of reference. Let us now analyze the aberration of light radiated by a single ”plane wave” emitter moving in the accelerated system. Before we go on to analyze an observations of a non-inertial observer, we should make one more remark about observations of an inertial observer. In the Chapter 4 we already emphasized that in the description of the aberration of light in the initial inertial frame of reference there are two choices of (four-dimensional) coordinatizations useful to consider:
(a) Non-standard (absolute time) coordinatization for a moving source
(b) Standard Lorentz coordinatization for a moving source
When the light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the light source at rest). This choice is usually the most convenient one from the viewpoint of connection to laboratory reality.
In fact, conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3 + 1) manifold. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time is absolute and we have no mixture of positions and time.
Let us, then, discuss (according to the Chapter 3) how to assign Lorentz coordinates in the case when a light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. In order to assign a Lorentz coordinate system in the inertial frame after the Galilean boost, one needs to perform the distant clock resynchronization $t^{\prime}=t-xv/c^{2}$. This new space-time coordinates in the lab frame are interpreted, mathematically, by saying that the metric is now diagonal and the speed of light from the moving source is isotropic and equal to $c$.
Now let us return to observations of a non-inertial observer. In the description of the aberration of light radiated by a single moving emitter in an accelerated frame of reference there are also two choices of coordinatizations:
(a) Non-standard (absolute time) coordinatization for an accelerated source
(b) Standard Lorentz coordinatization for an accelerated source
(a) When the system $S_{n}$ in the initial inertial lab frame $S$ is accelerated from the rest up to velocity $v$ along the $x$-axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. Acceleration with respect to the fixed stars have an effect on the propagation of light. The velocity of light from the source which is at rest in the accelerated frame $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have anisotropy along the $x_{n}$ axis (see Eq. (15)).
Figure 17: Aberration of light in an accelerated frame of reference $S_{n}$. Emitter is at rest with respect to the fixed stars. Wave fronts orientation and the group velocity presented in the Einstein synchronization. The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction. The aberration increment $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=-2v/c$, where $v$ is the velocity of the fixed stars in the accelerated frame $S_{n}$.
(b) When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction. Using such (Einstein time) synchronization procedure we actually describe the light from the source at rest (in the accelerated frame) using usual Maxwell’s equations.
@@ 5.8 A Single Moving Emitter in a Noninertial Frame of Reference
Above we considered the single emitter problem in the initial inertial frame of reference $S$. It is found that there is a deviation of the energy transport for the light radiated by the transversely moving emitter, which is nothing else but well-known (from textbooks) result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig. 1).
@@
We now consider the case when the emitter is at rest in the initial inertial frame $S$ (i.e. at rest with respect to the fixed stars) and the observer, which is at rest with respect to the accelerated frame of reference $S_{n}$ performs the direction of the energy transport measurement. It is important to emphasize that the aberration of light radiated by the single emitter is also dependent of the emitter motion in the accelerated system $S_{n}$. According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the emitter is at rest relative to the fixed stars and the observer started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame would jump by angle $-2v/c$. That is the aberration of light in the frame $S_{n}$ appears as shown in Fig 17.
We are going to discuss the asymmetry between inertial and accelerated frames. Now let us assume the absolute time coordinatization. The Langevin metric, together with initial condition, describes the aberration of light effect for the emitter at rest in the accelerated frame. The essential asymmetry is rooted in the electrodynamics equations. The laws of special relativity say that we can describe radiation of the moving with velocity $-v$ in the accelerated frame emitter in terms of the accelerated observer if we describe the electrodynamics by using metric
@@
This metric describes the electrodynamics of the moving light source with viewpoint of the accelerated observer measurements. From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x_{n}$ and time $t_{n}$ of the measured device at rest in the accelerated frame is the same as the measured result of this configuration at point $x_{n}+vt_{n}$ and time $t_{n}=0$. In the inertial frame we have a similar situation. In the Chapter 3 we described the electrodynamics of the moving emitter by using metric Eq. (11). This is the same way we used in the accelerated frame.
@@
So we begin to understand the basic machinery of the asymmetry between (initial) inertial and accelerated frames in the absolute time coordinatization: inertial observer describes the electrodynamics of the light source at rest using Minkowski metric. We substitute $x=x_{n}+vt,t_{n}=t$ into Minkowski metric to obtain Langevin metric Eq. 15 in the comoving frame. We may ask ourselves, what is the metric of the inertial observer with viewpoint of the accelerated observer? We transform coordinates $(t_{n},x_{n},y_{n},z_{n})$ that would be coordinates of an accelerated observer $S_{n}$ using the inverse Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Langevin metric Eq. 15 to obtain the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ in the initial inertial frame, as it must be. This metric describes the electrodynamics of the light source at rest in the inertial frame with viewpoint of the inertial observer measurements as viewing this of the accelerated observer. In particular this demonstrates the time dilation effect of a physical clock at rest in the accelerated frame. The slowly down of the accelerated clock does not depends on the choice of reference system, in which this effect is measured (see Chapter 9 for more details).
@@
There is still another way of solving the problem. The explanation consists in using the Langevin metric diagonalization procedure. The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for the emitter at rest in the accelerated frame and the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for the moving emitter. Now we see that there is a symmetry (reciprocity) in metrics (and consequently in electrodynamics equations) between the accelerated and inertial frames. This is consequence of the pseudo-Euclidean geometry of space-time. In fact, the difference between this metric and metric Eq. (11) is only the sign of velocity. Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions. The plane radiation wavefront of moving emitter (independently of its velocity) rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization. Of course in order to predict the result of aberration measurement the analysis can be done using directly the metric of the moving emitter in the absolute time coordinatization, Eq. (17). According to electrodynamics (see Section 5.4), the extra group velocity $-2v$ is introduced as a consequence of the crossed term $-4vdxdt$, the radiated beam of the moving emitter is propagated at the angle $-2v/c$.

### v10

@@ 5.5 A Resynchronization of the Accelerated Clocks
It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame $S_{n}$. The explanation consists of using a clock re-synchronization procedure. When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Using such a synchronization procedure we select a Lorentz coordinate system for the screen. In this synchronization, we describe the transmission through the aperture using Maxwell’s equations. The interval in the accelerated reference system $S_{n}$ will have the diagonal form Eq.(1) for the transmitted light beam.
@@
Figure 16: Aberration of light in the accelerated frame of reference $S_{n}$. The plane wavefront rotates in the accelerated frame after the accelerated clock re-synchronization. According to Maxwell’s electrodynamics, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
@@
The time $t^{\prime}_{n}$ under Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation. This time shift has the effect of rotation of the plane of simultaneity (that is emitter radiation wavefront) on the angle $-v/c$. As a consequence of this, the plane wavefront rotates in the accelerated frame after re-synchronization, Fig. 16. The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the light transmission description. Then, the transmitted (through the aperture) light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light: the two approaches give the same result. The choice between these two different clock synchronizations is a matter of pragmatics. By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear. But we can obtain a more consistent description of these phenomena.
@@
We had to restrict ourselves to finding the diagonalization procedure only for nonrelativistic velocities. The method can be easily generalized for an arbitrary parameter $v/c$. A full diagonalization can be obtained using transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Physical time $t^{\prime}_{n}$ determines the flow of time in a physical process in the accelerated frame $S_{n}$. The $dx^{\prime}_{n}$ is the physical distance between two points of 3-space.
@@ 5.6 Composition of Motions
Until now we have considered only the case in which motion is nonrelativistic in speed. Now we shall be more general and study the case where the motion is at an arbitrary velocity. We already know from our discussion in this section that metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the inverse Galilean boost. In this case, the metric of the inertial frame would pass into the Langevin metric of the accelerated frame. Let us consider the transformation of the direction of light propagation. The light (group) velocity transforms as a particle velocity. In this case, we use the Galileo addition theorem of velocities for transforming the speed of propagation of light to the accelerated frame. Let us assume that the accelerated frame moves with velocity $v$, then the angle of aberration $\theta_{a}$ is such that   $\tan\theta_{a}=-v/c$. If $v/c\ll 1$ the angle of aberration is small, being approximated by $\theta_{a}=-v/c$.
It is always possible to choose such variables, in which the metric of the accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at the transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. In fact, substituting $dx_{n}/dt_{n}=-v$ into this equation gives $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. The group velocity of the light beam along the fixed stars motion is still given by $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This is a modification of the aberration of light law in the Lorentz coordinatization. Suppose that an accelerated observer measures the light aberration. Since in the Lorentz coordinatization, the time and distance have direct physical meaning, the accelerated observer finds that the angle of aberration is such that $\sin\theta_{a}=-v/c$.
@@ 5.7 Discussion
Now let us return to metric diagonalization and try to get a better understanding of the relationship between the $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ coordinate systems. We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by the accelerated observer? Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in these frames should be related by the Lorentz transformation. At first sight, the problem is completely symmetrical and both frames are equivalent. Above we demonstrated that there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history.
@@
Where does the asymmetry come from? The electrodynamics equation needs to be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical direction $z$. On the other hand, we know that when we are dealing with Lorentz coordinatization we have an effect on the wavefront orientation in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
@@
This is purely a kinematic effect, involving no forces. Many people will, we think, find this disturbing. We discuss the key to this paradox in great detail in Section 12. It is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. Without proof, we may state the results. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the system $S_{n}$ as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process. The rate of a clock depends on the pseudo-gravitational potential at the place where the clock is situated. The time-offset relation can be interpreted as the accumulated time difference between spatially separated clocks. Suppose (for simplicity) that there is a constant acceleration $g$ along the $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. The clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis, we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks, we have $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift has the effect of rotation the plane of simultaneity on the angle $-v/c$.
5.8 A Single Moving Emitter in an Accelerated Frame of Reference
@@
Above in Section 4, we considered the single emitter problem in the initial inertial frame of reference $S$. It is found that there is a deviation of the energy transport for the light radiated by the accelerated emitter, which is nothing else but a well-known result: there is the aberration of light from the tangentially moving emitter in the inertial frame of reference, Fig. 1. We demonstrated that there could be two (covariant and no-covariant) approaches to the analysis of the aberration of light radiated by a single moving emitter.
Let us now analyze the aberration of light radiated by a single ”plane wave” emitter moving in the accelerated system. In this case, there are also two choices of coordinatizations:
(a) When the system $S_{n}$ in the initial inertial lab frame $S$ is accelerated from the rest up to velocity $v$ along the $x$-axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. Acceleration with respect to the fixed stars affects the propagation of light. The velocity of light from the source which is at rest in the accelerated frame $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have anisotropy along the $x_{n}$ axis (see Eq. (15)).
Figure 17: Aberration of light in an accelerated frame of reference $S_{n}$. The emitter is at rest with respect to the fixed stars. Wavefront orientation and the group velocity are presented in the Einstein synchronization.
(b) When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all directions. Using such (Einstein time) synchronization procedure we describe the light from the source at rest (in the accelerated frame) using usual Maxwell’s equations.
@@
We now consider the case when the emitter is at rest in the initial inertial frame $S$ (i.e. at rest with respect to the fixed stars) and the observer, which is at rest with respect to the accelerated frame of reference $S_{n}$ performs the direction of the energy transport measurement. It is important to emphasize that the aberration of light radiated by the single emitter is also dependent on the emitter motion in the accelerated system $S_{n}$. According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the emitter is at rest relative to the fixed stars and the observer started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame would jump by angle $-2v/c$. That is the aberration of light in the frame $S_{n}$ appears as shown in Fig 17.
Let us describe the radiation of the moving with velocity $-v$ in the accelerated frame emitter in terms of the accelerated observer. Acceleration with respect to the accelerated frame affects the propagation of light. Let us assume the absolute time coordinatization. Using the Galilean boost: we substitute $x_{n}\to x_{n}+vt_{n}$, while leaving time unchanged into the Langevin metric to obtain
@@
This metric describes the electrodynamics of the light source at rest in the inertial frame with the viewpoint of the accelerated observer measurements. From the formula itself, it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x_{n}$ and time $t_{n}$ of the measured device at rest in the accelerated frame is the same as the measured result of this configuration at point $x_{n}+vt_{n}$ and time $t_{n}=0$. In the inertial frame, we have a similar situation. In Chapter 3 we described the electrodynamics of the moving emitter by using metric Eq. (11). Within a single accelerated frame, in the inertial stage, the symmetry principle (i.e. equivalence of the active and passive Galilean boost) is satisfied.
@@
Let us now discuss the basic machinery of the asymmetry between (initial) inertial and accelerated frames in absolute time coordinatization: inertial observer describes the electrodynamics of the light source at rest using the Minkowski metric. We substitute $x=x_{n}+vt,t_{n}=t$ into the Minkowski metric to obtain the Langevin metric Eq. 15 in the comoving frame. We may ask ourselves, what is the metric of the inertial observer with the viewpoint of the accelerated observer? We transform coordinates $(t_{n},x_{n},y_{n},z_{n})$ that would be coordinates of an accelerated observer $S_{n}$ using the inverse Galilean transformation: we substitute $x_{n}=x-vt$ while leaving time unchanged $t_{n}=t$ into the Langevin metric Eq. 15 to obtain the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ in the initial inertial frame, as it must be. This metric describes the electrodynamics of the light source at rest in the inertial frame with the viewpoint of the inertial observer measurements as viewing this of the accelerated observer. In particular, this demonstrates the time dilation effect of a physical clock at rest in the accelerated frame. The slowdown of the accelerated clock does not depend on the choice of the reference system, in which this effect is measured (see Chapter 12 for more details).
@@
There is still another way of solving the problem. The explanation consists of using the Langevin metric diagonalization procedure. The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for the emitter at rest in the accelerated frame and the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for the moving emitter. Now we see that there is a symmetry (reciprocity) in metrics (and consequently in electrodynamics equations) between the accelerated and inertial frames. This is a consequence of the pseudo-Euclidean geometry of space-time. The difference between this metric and metric Eq. (11) is only the sign of velocity. Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions. The plane radiation wavefront of the moving emitter (independently of its velocity) rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization. Of course, to predict the result of aberration measurement the analysis can be done using directly the metric of the moving emitter in the absolute time coordinatization, Eq. (18). According to electrodynamics, the group velocity $-2v$ is introduced as a consequence of the crossed term $-4vdxdt$.